In a game of football outdoors on a cold day, a player will begin to feel exhausted after using approximately 8.0 × 105 J of internal energy. (a) One player, dressed too lightly for the weather, has to leave the game after losing 6.4 × 105 J of heat. How much work has he done? (b) Another player, wearing clothes that offer better protection against heat loss, is able to remain in the game long enough to do 2.8 × 105 J of work. What is the magnitude of the heat that he has lost?
Answer #1

Solution -

This question can be solved by the use of first law of thermodynamics -

\Delta U = Q - W

In first case

  • change in internal energy is = - 8.0 × 105 J ( - ve sign is used to tell that the internal energy of body is used hence decreasing )
  • Heat loss = - 6.4 × 105 J (-ve sign shows heat loss from the system )

Putting into the equation -

- 8.0 × 105 J =  - 6.4 × 105 J - W

W = 1.6  × 105 J ( This is the work done by the player )

In second case -

  • W = 2.8 × 105 J ( Positive because the work is being done by the system )
  • change in internal energy is = - 8.0 × 105 J ( - ve sign is used to tell that the internal energy of body is used hence decreasing )

Putting into the equation -

- 8.0 × 105 J =  Q - 2.8 × 105 J

Q = - 5.2 × 105 J ( This is the amount of heat loss by the player )  

If you finds the answer conceptually good and correct , please rate good and comment your feedback .

Thank you .